OpenStudy (anonymous):

2x^2-4x-18=0 Complete the square

OpenStudy (mathstudent55):

1. Add 18 to both sides to get the -18 term on the right side.

OpenStudy (mathstudent55):

2. You must have x^2 as the first term, so divide both sides by 2. Do these two steps and show what you got.

OpenStudy (luigi0210):

Go!

OpenStudy (anonymous):

Simplifying 2x2 + 4x + -18 = 0 Reorder the terms: -18 + 4x + 2x2 = 0 Solving -18 + 4x + 2x2 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '2'. 2(-9 + 2x + x2) = 0 Ignore the factor 2. Subproblem 1 Set the factor '(-9 + 2x + x2)' equal to zero and attempt to solve: Simplifying -9 + 2x + x2 = 0 Solving -9 + 2x + x2 = 0 Begin completing the square. Move the constant term to the right: Add '9' to each side of the equation. -9 + 2x + 9 + x2 = 0 + 9 Reorder the terms: -9 + 9 + 2x + x2 = 0 + 9 Combine like terms: -9 + 9 = 0 0 + 2x + x2 = 0 + 9 2x + x2 = 0 + 9 Combine like terms: 0 + 9 = 9 2x + x2 = 9 The x term is 2x. Take half its coefficient (1). Square it (1) and add it to both sides. Add '1' to each side of the equation. 2x + 1 + x2 = 9 + 1 Reorder the terms: 1 + 2x + x2 = 9 + 1 Combine like terms: 9 + 1 = 10 1 + 2x + x2 = 10 Factor a perfect square on the left side: (x + 1)(x + 1) = 10 Calculate the square root of the right side: 3.16227766 Break this problem into two subproblems by setting (x + 1) equal to 3.16227766 and -3.16227766. Subproblem 1 x + 1 = 3.16227766 Simplifying x + 1 = 3.16227766 Reorder the terms: 1 + x = 3.16227766 Solving 1 + x = 3.16227766 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + x = 3.16227766 + -1 Combine like terms: 1 + -1 = 0 0 + x = 3.16227766 + -1 x = 3.16227766 + -1 Combine like terms: 3.16227766 + -1 = 2.16227766 x = 2.16227766 Simplifying x = 2.16227766 Subproblem 2 x + 1 = -3.16227766 Simplifying x + 1 = -3.16227766 Reorder the terms: 1 + x = -3.16227766 Solving 1 + x = -3.16227766 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + x = -3.16227766 + -1 Combine like terms: 1 + -1 = 0 0 + x = -3.16227766 + -1 x = -3.16227766 + -1 Combine like terms: -3.16227766 + -1 = -4.16227766 x = -4.16227766 Simplifying x = -4.16227766 Solution The solution to the problem is based on the solutions from the subproblems. x = {2.16227766, -4.16227766} Solution x = {2.16227766, -4.16227766}